Circular No. 6971

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cp IN MCG +9-21-11 A. V. Filippenko and A. J. Barth, University of California at Berkeley,

report that a CCD spectrum obtained on July 17 UT with the Shane 3m reflector at Lick Observatory confirms that SN 1998cp (cf. IAUC 6954) is indeed a supernova. However, the spectral classification is uncertain; further processing of the data is required.

XTE J2123–058 J. Casares, M. Serra-Ricart, C. Zurita, A. Gomez, and D. Alcalde, In-

stituto de Astrofisica de Canarias; and P. Charles, University of Oxford, report the discovery of optical eclipses in the x-ray transient XTE J2123– 058 (IAUC 6955): “R-band images were obtained between July 2 and 12 with the European Space Agency’s Optical Ground Station 1-m telescope at the Observatorio del Teide. Our light curves show triangular-shaped minima with a central depth of 0.65 mag, repeating each night, indicating a periodicity of roughly 1 day. The dips, which we interpret as eclipses of the accretion disk by the companion star, have a total duration of 0.16 day. Preliminary analysis, using phase-dispersion minimization, yields a strong peak at 1.99 days. After phase folding on this period and submultiples of it, we favor an orbital period of 0.993 or 0.993/2 day. A provisional ephemeris for the time of mid-eclipse is HJD 24451001.685(2) + 0.993(2)E. Observations at other wavelengths (especially in x-rays) to monitor the decline are strongly encouraged. Optical monitoring from widely-distributed sites is particularly important to disentangle 1-day aliases and search for the secondary minimum.”

J. Homan, M. van der Klis, J. van Paradijs, and M. Mendez, Astronomical Institute ‘Anton Pannekoek’, University of Amsterdam, communicate: “We report the discovery, using RXTE PCA data from July 14.4 to 14.8 UT, of two high-frequency quasiperiodic oscillations (QPOs) in the persistent x-ray flux of the newly discovered x-ray transient XTE J2123–058 (IAUC 6955). The QPOs have frequencies of 855 ± 5 and 1123 ± 7 Hz, widths of 33.5 ± 11.8 and 51.5 ± 15.1 Hz, and rms amplitudes of 6.3 ± 0.8 and 7.4 ± 0.9 percent (2–11 keV). Both QPOs get stronger with photon energy and are detected at count rates below ∼ 640 counts/s (2–60 keV, including background), but undetected above. Two x-ray bursts were examined for oscillations, but none were found. We recommend further RXTE observations when the PCA level drops below ∼ 640 counts/s.”

1998 July 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green